# Generalized Kummer surfaces over finite fields

### v3

@@ Abstract.
In this paper, we prove a refinement of the Katsura theorem on finite groups acting on abelian surfaces such that the quotient is birational to a $K3$ surface. As an application, we compute traces of Frobenius on the Neron–Severi groups of supersingular generalized Kummer surfaces over finite fields.
@@
Finite field, abelian variety, Kummer surface
@@ 1. Introduction
Throughout this paper $k$ is a perfect field of characteristic $p$, and ${\bar{k}}$ is an algebraic closure of $k$. Let $A$ be an abelian surface over $k$ with an action of a finite group $G$. In this paper, we study only actions such that $G$ preserves the group law on $A$; in other words, there is a homomorphism of the groups $G\to\operatorname{End}(A)^{*}$, where for a ring $R$ we denote by $R^{*}$ the multiplicative group of invertible elements. If a minimal smooth model $X(A,G)$ of the quotient $A/G$ is a $K3$ surface, we say that $X(A,G)$ is a generalized Kummer surface. In [Ry12] we classified zeta functions of Kummer surfaces $X(A,{\mathbb{Z}}/2{\mathbb{Z}})$ over finite fields of characteristic $p>2$. This paper is a step towards such a classification for generalized Kummer surfaces.
For an algebraic variety $X$ over $k$ we denote by $\bar{X}$ the base change of $X$ to ${\bar{k}}$. A $K3$ surface $X$ is called Shioda supersingular if the rank of $\operatorname{NS}(\bar{X})$ is $22$. A Shioda supersingular $K3$ surface is supersingular, and a supersingular Kummer surface is Shioda supersingular [Ar]. For a Shioda supersingular $K3$ surface we have the natural isomorphism for any prime $\ell\neq p$:
()( ∗ ) $\operatorname{NS}(\bar{X})\otimes{\mathbb{Q}}_{\ell}\to H^{2}(\bar{X},{\mathbb{Q}}_{\ell}).$
It follows that, if the base field $k$ is finite, the zeta function of a Shioda supersingular $K3$ surface $X$ is uniquely determined by the Frobenius action on the Neron-Severi group $\operatorname{NS}(\bar{X})$. This observation is a link between the arithmetic and geometry of supersingular $K3$ surfaces.
Since odd cohomology groups of $X$ are trivial, the isomorphism ()(*)( ∗ ) leads to the formula for the number of points on $X$ over $k$:
()( ∗ ∗ ) $|X(k)|=1+q\operatorname{Tr}_{X}+q^{2},$
@@
Our main motivation comes from the analogy with cubic surfaces in ${\mathbb{P}}^{3}$. For a cubic surface $X$ we have the same formula ()(**)( ∗ ∗ ) for the number of points. In [Serre12] Serre asked which values of $\operatorname{Tr}_{X}$ can arise for smooth cubic surfaces in ${\mathbb{P}}^{3}$ over ${\mathbb{F}}_{q}$. The answer was recently obtained in [BFL]. A more general question on zeta functions of cubic surfaces was treated in [RT16, T20, LT20]. A direct generalization of Serre’s question to quartics in ${\mathbb{P}}^{3}$ seems very difficult, but we can try to generalize the question as follows. Since a quartic in ${\mathbb{P}}^{3}$ with reasonable singularities is a $K3$ surface, we can study traces of Frobenius on Neron–Severi groups of $K3$ surfaces. The isomorphism ()(*)( ∗ ) holds for any smooth cubic surface, but not for any $K3$ surface; therefore, we focus on Shioda supersingular $K3$ surfaces. In this paper, we compute Frobenius traces for generalized Kummer surfaces only, but it is natural to pose the question for supersingular $K3$ surfaces as well.
We heavily use results from the pioneering paper of Katsura on generalized Kummer surfaces over fields of positive characteristics [Ka87]. We reformulate some of his results here in a more convenient way.
Definition 1.1.
We say that the action of a finite group $G$ on an abelian variety $A$ is rigid if the representation of $G$ in $V_{\ell}(A)$ is without fixed points, i.e., for any $g\in G$ of order $r$ the eigenvalues of the action of $g$ on $V_{\ell}(A)$ are primitive roots of unity of degree $r$.
@@
If an action of $G$ on $A$ is rigid, then it is faithful. Moreover, if $g\in G$ is an element of order $2$, then $g$ acts on $A$ as $-1$; thus there is at most one element of order $2$ in $G$.
@@
Katsura classified finite groups $G$ such that there exists a finite field $k$ of characteristic $p>5$ and an abelian surface $A$ over $k$ with an action of $G$ such that the quotient is birational to a $K3$ surface.
@@
[Ka87, Theorem 3.7] Let $k$ be a field of characteristic $p>5$. Let $A$ be an abelian surface over $k$ with an action of a finite group $G$ such that the quotient is birational to a $K3$ surface. If all elements of $G$ fix the group law of $A$, then $G$ is one of the following groups:
@@
Katsura showed that any group from Theorem 1.4 occurs over some finite field of characteristic $p>5$. As far as author knows, there is no classification of such groups over a given finite field. In section 6 we prove the following statement.
@@
    $G$ is $Q_{8},Q_{12},\operatorname{SL}_{2}({\mathbb{F}}_{3})$, or a cyclic group of order $2,3,4$, or $6$;
@@
    if $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then $A$ is supersingular, ${\mathbb{F}}_{p^{2}}\subset k$, and $p\not\equiv\pm 1\bmod n$.
To prove the theorem, we use that there is a rigid action of a finite group $G$ on an abelian variety $A$ if and only if there is a homomorphism from a certain semisimple algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ to the endomorphism algebra $\operatorname{End}(A)\otimes{\mathbb{Q}}$ of $A$ (see Section 3). The algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ is the rigid quotient of the group algebra ${\mathbb{Q}}[G]$. In Section 4 we obtain a classification of groups $G$ with a rigid action on an abelian surface. This result can be extracted from the results of Hwang [Hw], but we feel that in this particular case it is easier to give an independent proof.
@@
There exist supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$, where $q$ is an even power of $p$, with the following traces: $22,18,14,10,8,6,4,2,0$
There exist supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$, where $q$ is an odd power of $p$, with the following traces: $20,18,14,10,8,6,2,0$
@@
The author thanks Yuri Zarhin and Matthias Schütt for interesting discussions and WonTae Hwang for his remarks on the paper. The author is a Simons-IUM contest winner and thanks its sponsors and jury.
@@ 2.1. Notation
[…]
2.2. Central simple algebras over number fields
Recall some well-known facts on central simple algebras over number fields [MO, Section 28]. Let $L/K$ be an extension of number fields, and let ${\mathbb{H}}$ be a simple algebra with center $K$ such that $L\otimes{\mathbb{H}}$ is a matrix algebra over $L$, i.e., ${\mathbb{H}}$ represents an element of the Brauer group $\operatorname{Br}(L/K)$. Let $v$ be a prime of $K$. We denote the local invariant of ${\mathbb{H}}$ in $v$ by $\operatorname{inv}_{v}{\mathbb{H}}\in{\mathbb{Q}}/{\mathbb{Z}}$. A simple algebra over a number field $K$ is uniquely determined by its local invariants in the places of $K$ [MO, Remark 32.12].
@@ 2.2. Central simple algebras over number fields
Let ${\mathbb{H}}$ be a central simple algebra over $K$, and let $v$ be a prime of $K$. Let $L$ be an extension of $K$. Choose a prime $w$ of $L$ over $v$. Then ${\mathbb{H}}_{L}$ is a central simple algebra over $L$ with local invariant $[L_{w}:K_{v}]\operatorname{inv}_{v}{\mathbb{H}}$ in $w$.
@@
Let $K$ be a totally real extension of ${\mathbb{Q}}$ of even degree. We denote by ${\mathbb{H}}_{\infty}(K)$ the quaternion algebra with center $K$, and non-trivial invariants exactly in the real places of $K$. We denote by ${\mathbb{H}}_{p}$ the quaternion algebra with center ${\mathbb{Q}}$, and non-trivial invariants in the real place and in $p$.
Theorem 2.2.
@@
It follows from [MO, Theorem 28.5]. ∎
@@
    Let $K$ be a real quadratic extension of ${\mathbb{Q}}$. There is a homomorphism of ${\mathbb{H}}_{p}$ to ${\mathbb{H}}_{\infty}(K)$ if and only if $p$ is not split in $K$.
@@
    Let $L$ be an imaginary quadratic extension of ${\mathbb{Q}}$. There is a homomorphism of ${\mathbb{H}}_{p}$ to $M(2,L)$ if and only if $p$ is not split in $L$.
@@
    If $p$ is not split in $K$, then, according to Theorems 2.1 and 2.2, we have a homomorphism
@@
    Assume that $p$ is split in $K$, and that there is a homomorphism ${\mathbb{H}}_{p}\to{\mathbb{H}}_{\infty}(K)$. Then we get a homomorphism of the simple algebra ${\mathbb{H}}_{p}\otimes{\mathbb{Q}}_{p}$ to the sum of two matrix algebras
@@
    There is a homomorphism of ${\mathbb{H}}_{p}$ to $M(2,L)$ if and only if there is a homomorphism in the opposite direction of the centralizers of these algebras in $M(4,{\mathbb{Q}})$, i.e., $L\to{\mathbb{H}}_{p}$. According to Theorems 2.1 and 2.2, the last homomorphism exists if and only if $p$ is not split in $L$.
@@ 2.3. Abelian varieties over finite fields
Let $A$ be an abelian variety over $k$. The endomorphism ring of an abelian variety $\operatorname{End}(A)$ is finitely generated and torsion-free as ${\mathbb{Z}}$-module. Let
@@
An element $\varphi\in\operatorname{End}(A)$ is called an isogeny if $\varphi$ is finite and surjective. For example, multiplication by $m\in{\mathbb{Z}}$ is an isogeny $[m]:A\to A$. Denote the kernel of $[m]$ by $A[m]$. This is a finite group scheme over $k$ of order $m^{2\dim A}$ [Mil08, Remark 7.3].
An abelian variety $A$ is simple if it does not contain nontrivial abelian subvarieties. Any abelian variety $A$ over $k$ is isogenous to a product of simple abelian varieties:
@@
where $A_{i}$ are simple. This decomposition corresponds to a decomposition of $\operatorname{End}^{\circ}(A)$ into a product of simple algebras $\operatorname{Mat}(r_{i},\operatorname{End}^{\circ}(A_{i}))$. In particular, $\operatorname{End}^{\circ}(A)$ is a semisimple ${\mathbb{Q}}$-algebra [Mum70, IV.19, Corollaries 1 and 2].
@@
be the $\ell$-th Tate module of $A$, and the corresponding vector space over ${\mathbb{Q}}_{\ell}$. It is known that $T_{\ell}(A)$ is a free ${\mathbb{Z}}_{\ell}$-module of rank $2\dim A$.
Let $k$ be a finite field ${\mathbb{F}}_{q}$. The Frobenius endomorphism $F_{A}:A\to A$ over $k$ is the morphism that is trivial on the topological space and raises functions to their $q$-th powers. The endomorphism $F_{A}$ acts on the Tate module by a semisimple linear transformation, which we denote by $F:T_{\ell}(A)\to T_{\ell}(A)$. The characteristic polynomial
@@
is called the Weil polynomial of $A$. It is a monic polynomial of degree $2\dim A$ with rational integer coefficients independent of the choice of $\ell$ (see [Mum70, Page 180, Theorem 4], and [Mum70, Page 205]).
The Frobenius endomorphism generates the center $K_{A}\subset\operatorname{End}^{\circ}(A)$. Tate proved that the isogeny class of an abelian variety over $k$ is determined by its characteristic polynomial, that is $f_{A}(t)=f_{B}(t)$ if and only if $A$ is isogenous to $B$ [Ta66]. Moreover, for any $\ell\neq p$ we have
@@
If $f:A\to B$ is an isogeny, then $T_{\ell}(f):T_{\ell}(A)\to T_{\ell}(B)$ is injective, and the preimage $T=T_{\ell}(f)^{-1}(T_{\ell}(B))$ is an $F$-invariant submodule of $V_{\ell}(A)$. Conversely, if $T\subset V_{\ell}(A)$ is $F$-invariant ${\mathbb{Z}}_{\ell}$-submodule of finite rank such that $T_{\ell}(A)\subset T$, then there exists an abelian variety $B$ defined over $k$ and an $\ell$-isogeny $f:A\to B$ such that $T_{\ell}(f)$ induces an isomorphism $T\cong T_{\ell}(B)$.∎
Assume that $A$ is simple. In this case $K_{A}$ is a field. Let $L_{A}$ be the Galois envelope of $K_{A}$ over ${\mathbb{Q}}$, and let $P$ be the set of primes of $L_{A}$ over $p$. The slopes of $A$ are defined as the set
@@
where $v_{\mathfrak{p}}$ is the normalized valuation. We say that $A$ is supersingular if all slopes of $A$ are equal to $1/2$.
@@
Let $L$ be a Galois extension of ${\mathbb{Q}}$. Assume that $\sqrt{q}\not\in{\mathbb{Z}}$, and $p$ is unramified in $L$. Then there is no supersingular abelian variety $A$ such that $L_{A}=L$.
@@
Let $A$ be an abelian variety such that $L_{A}=L$. Then the slopes of $A$ are the fractions ${v_{\mathfrak{p}}(F_{A})}/{v_{\mathfrak{p}}(q)}$ with odd denominator. ∎
The abelian variety is ordinary, if the set of slopes of $A$ is $\{0,1\}$. It is known that $A$ is ordinary if and only if the order of the group $A[p]({\bar{k}})$ is equal to $p^{\dim A}$. A simple abelian surface $A$ is called mixed, if the set of slopes of $A$ is $\{0,1/2,1\}$.
In dimensions $1$ and $2$ Weil polynomials can be explicitly classified. We present this classification in part. For more details, see [MN02] and [Wa69].
@@
[Wa69, Ru90, MN02] Let $A$ be a simple abelian surface over ${\mathbb{F}}_{q}$, where $q=p^{n}$. Then $f_{A}(t)=P_{A}(t)^{e}$, where $P_{A}\in{\mathbb{Z}}[t]$ is irreducible.
@@
    If $A$ is ordinary or mixed, then $e=1$, and
@@
    If $A$ is supersingular, then we have the following possibilities:
@@
        $(0,-q)$, $n$ is even, and $p\not\equiv 1\bmod 12$;
@@
        (±$(\pm\sqrt{q},q)$, $n$ is even, and $p\not\equiv 1\bmod 5$;

### v4

@@ Abstract.
We prove a refinement of the Katsura theorem on finite group actions on abelian surfaces such that the quotient is birational to a $K3$ surface. As an application, we compute traces of Frobenius on the Neron–Severi groups of supersingular generalized Kummer surfaces over finite fields.
@@
Finite field; abelian variety; Kummer surface; generalized Kummer surface
@@ 1. Introduction
Throughout this paper $k$ is a perfect field of characteristic $p$, and $\bar{k}$} is an algebraic closure of $k$. Let AA be an abelian surface over $k$ with an action of a finite group $G$. In this paper, we study only actions such that $G$ preserves the group law on AA; in other words, there is a group homomorphism $G\to\operatorname{End}(A)^{*}$, where for a ring $R$ we denote by $R^{*}$ the multiplicative group of invertible elements. If a minimal smooth model $X(A,G)$ of the quotient $A/G$ is a $K3$ surface, we say that $X(A,G)$ is a generalized Kummer surface. In [Ry12] we classified the zeta functions of Kummer surfaces $X(A,{\mathbb{Z}}/2{\mathbb{Z}})$ over finite fields of characteristic $p>2$. This paper is a step towards such a classification for generalized Kummer surfaces.
For an algebraic variety $X$ over $k$ we denote by $\bar{X}$ the base change of $X$ to $\bar{k}$}. A $K3$ surface $X$ is called Shioda supersingular if the rank of $\operatorname{NS}(\bar{X})$ is $22$. A Shioda supersingular K3 surface is supersingular, and a supersingular Kummer surface is Shioda supersingular [Ar74]. More generally, the Tate conjecture is proved for $K3$ surfaces over finitely generated fields of odd characteristic [MP15]; therefore, a $K3$ surface over such a field is Shioda supersingular if and only if it is supersingular.
For a Shioda supersingular $K3$ surface $X$, we have the natural isomorphism for any prime $\ell\neq p$:
() $\operatorname{NS}(\bar{X})\otimes{\mathbb{Q}}_{\ell}\to H^{2}(\bar{X},{\mathbb{Q}}_{\ell}).$
We see that if the base field $k$ is finite, the zeta function of a Shioda supersingular $K3$ surface $X$ is uniquely determined by the Frobenius action on the Neron-Severi group $\operatorname{NS}(\bar{X})$. This observation is a link between the arithmetic and geometry of a supersingular $K3$ surface.
Since odd cohomology groups of $X$ are trivial, the isomorphism ()(*) leads to the formula for the number of points on $X$ over $k$:
() $|X(k)|=1+q\operatorname{Tr}_{X}+q^{2},$
@@
Our main motivation comes from the analogy with the cubic surfaces in ${\mathbb{P}}^{3}$. For a cubic surface $X$ we have the same formula ()(**) for the number of points. In [Serre12] Serre asked which values of $\operatorname{Tr}_{X}$ can arise for smooth cubic surfaces in ${\mathbb{P}}^{3}$ over ${\mathbb{F}}_{q}$. The answer was recently obtained in [BFL]. A more general question about the zeta functions of cubic surfaces was treated in [RT16, T20, LT20]. A direct generalization of Serre’s question to quartics in ${\mathbb{P}}^{3}$ seems very difficult, but we can try to generalize the question as follows. Since a quartic in ${\mathbb{P}}^{3}$ with reasonable singularities is a $K3$ surface, we can study traces of Frobenius on Neron–Severi groups of $K3$ surfaces. The isomorphism ()(*) holds for any smooth cubic surface, but not for any $K3$ surface; therefore, we focus on Shioda supersingular $K3$ surfaces. In this paper we compute Frobenius traces for generalized Kummer surfaces only, but it is natural to pose the question for supersingular $K3$ surfaces as well.
The paper is organized as follows. In Section 3 we study rigid group actions on abelian varieties.
Definition $1.$
We say that the action of a finite group $G$ on an abelian variety AA is rigid if the representation of $G$ in $V_{\ell}(A)$ is without fixed points, i.e., for any $g\in G$ of order $r$ the eigenvalues of the action of $g$ on $V_{\ell}(A)$ are primitive roots of unity of degree $r$.
@@
If an action of $G$ on AA is rigid, then it is faithful. Moreover, if $g\in G$ is an element of order 22, then $g$ acts on AA as 1-1; thus there is at most one element of order 22 in $G$.
${\mathbb{Q}}[G]=\oplus_{V}{\mathbb{H}}_{V}.$
${\mathbb{Q}}[G]^{\mathrm{rig}}:=\oplus_{V\text{ is WFP}}{\mathbb{H}}_{V}.$
The following result is an immediate consequence of Theorem 3.5.
Theorem 1.3.
There exists an abelian variety with a rigid action of $G$ in the isogeny class of an abelian variety AA, if and only if there exists a homomorphism of $\mathbb{Q}$}-algebras ${\mathbb{Q}}[G]^{\mathrm{rig}}\to\operatorname{End}^{\circ}(A).$
In Section 4 we use this result to obtain a classification of finite groups $G$ with a rigid action on an abelian surface. This result can be extracted from the results of Hwang [Hw21], but we feel that in this particular case it is easier to give an independent proof. Section 5 is devoted to singularities of $A/G$.
In Section 6 we heavily use results from the pioneering paper of Katsura on generalized Kummer surfaces over a field of positive characteristic [Ka87]. Let us reformulate some of his results here in a more convenient way.
@@
Katsura classified finite groups $G$ such that there exists a finite field $k$ of characteristic $p>5$ and an abelian surface AA over $k$ with an action of $G$ such that the quotient is birational to a $K3$ surface.
@@
[Ka87, Theorem 3.7] Let $k$ be a field of characteristic $p>5$. Let AA be an abelian surface over $k$ with an action of a finite group $G$ such that the quotient is birational to a $K3$ surface. If all elements of $G$ fix the group law of AA, then $G$ is one of the following groups:
@@
Katsura showed that any group from Theorem 1.5 occurs over some finite field of characteristic $p>5$. As far as the author knows, there is no classification of such groups over a given finite field. In Section 6 we prove the following statement.
@@
    $G$ is $Q_{8},Q_{12},\operatorname{SL}_{2}({\mathbb{F}}_{3})$, or a cyclic group of order $2,3,4$, or 66;
@@
    if $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then AA is supersingular, ${\mathbb{F}}_{p^{2}}\subset k$, and $p\not\equiv\pm 1\bmod n$.
@@
There exist supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$, where $q$ is an even power of $p$, with the following traces: $22,18,14,10,8,6,4,2,0.$
There exist supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$, where $q$ is an odd power of $p$, with the following traces: $20,18,14,10,8,6,2,0.$
@@
Acknowledgements
The work was supported by the Israel Science Foundation, grant No. 1405/22 I would like to thank Yuri Zarhin for interesting discussions and Matthias Schütt for email exchanges. I also thank WonTae Hwang for his remarks on the paper. I am a Simons-IUM contest winner, and I am grateful to its sponsors and jury.
@@ 2.1. Notation
$\zeta_{n}:$ $\text{ a primitive root of unity of order }n;$
$\Phi_{r}:$ $\text{ the }r\text{-th cyclotomic polynomial};$
[…]
$M(r,{\mathbb{H}}):$ square matrices over an algebra $\text${ square matrices over an algebra $\mathbb{H}$};
@@
${{\mathbb{Q}}[G]}^{\mathrm{rig}}:$ $\text{ the rigid quotient algebra of }{\mathbb{Q}}[G].$
$2.$ Central simple algebras over number fields
Recall some well-known facts on central simple algebras over number fields [MO, Section 28]. Let $L/K$ be an extension of number fields, and let $\mathbb{H}$} be a simple algebra with center $K$ such that $L\otimes{\mathbb{H}}$ is a matrix algebra over $L$, i.e., $\mathbb{H}$} represents an element of the Brauer group $\operatorname{Br}(L/K)$. Let $v$ be a prime of $K$. We denote the local invariant of $\mathbb{H}$} at $v$ by $\operatorname{inv}_{v}{\mathbb{H}}\in{\mathbb{Q}}/{\mathbb{Z}}$. A simple algebra over a number field $K$ is uniquely determined by its local invariants at the places of $K$ [MO, Remark 32.12].
@@
Let $\mathbb{H}$} be a central simple algebra over $K$, and let $v$ be a prime of $K$. Let $L$ be an extension of $K$. Choose a prime $w$ of $L$ over $v$. Then ${\mathbb{H}}_{L}$ is a central simple algebra over $L$ with local invariant $L_{w}:K_{v}]\operatorname{inv}_{v}{\mathbb{H}}$ at $w$.
@@
Let $K$ be a totally real extension of $\mathbb{Q}$} of even degree. We denote by ${\mathbb{H}}_{\infty}(K)$ the quaternion algebra with center $K$, and non-trivial invariants exactly at the real places of $K$. We denote by ${\mathbb{H}}_{p}$ the quaternion algebra with center $\mathbb{Q}$}, and non-trivial invariants at the real place and at pp.
Theorem $2.$
@@
This follows from [MO, Theorem 28.5]. ∎
@@
    Let $K$ be a real quadratic extension of $\mathbb{Q}$}. There is a homomorphism of ${\mathbb{H}}_{p}$ to ${\mathbb{H}}_{\infty}(K)$ if and only if $p$ does not split in $K$.
@@
    Let $L$ be an imaginary quadratic extension of $\mathbb{Q}$}. There is a homomorphism of ${\mathbb{H}}_{p}$ to $M(2,L)$ if and only if $p$ does not split in $L$.
@@
    If $p$ does not split in $K$, then, according to Theorems 2.1 and 2.2, we have a homomorphism
@@
    Assume that $p$ splits in $K$, and that there is a homomorphism ${\mathbb{H}}_{p}\to{\mathbb{H}}_{\infty}(K)$. Then we get a homomorphism of the simple algebra ${\mathbb{H}}_{p}\otimes{\mathbb{Q}}_{p}$ to the sum of two matrix algebras
@@
    There is a homomorphism of ${\mathbb{H}}_{p}$ to $M(2,L)$ if and only if there is a homomorphism in the opposite direction of the centralizers of these algebras in $M(4,{\mathbb{Q}})$, that is, $L\to{\mathbb{H}}_{p}$. According to Theorems 2.1 and 2.2, the last homomorphism exists if and only if $p$ does not split in $L$.
@@ 2.3. Abelian varieties over finite fields
Let AA be an abelian variety over $k$. The endomorphism ring of an abelian variety $\operatorname{End}(A)$ is finitely generated and torsion-free as $\mathbb{Z}$}-module. Let
@@
An element $\varphi\in\operatorname{End}(A)$ is called an isogeny if $\varphi$ is finite and surjective. For example, multiplication by $m\in{\mathbb{Z}}$ is an isogeny $m]:A\to A$. Denote the kernel of $[m]$ by $A[m]$. This is a finite group scheme over $k$ of order $m^{2\dim A}$ [Mil08, Remark 7.3].
An abelian variety AA is simple if it does not contain non-trivial abelian subvarieties. Any abelian variety AA over $k$ is isogenous to a product of simple abelian varieties:
@@
where $A_{i}$ are simple. This decomposition corresponds to a decomposition of $\operatorname{End}^{\circ}(A)$ into a product of simple algebras $\operatorname{Mat}(r_{i},\operatorname{End}^{\circ}(A_{i}))$. In particular, $\operatorname{End}^{\circ}(A)$ is a semi-simple $\mathbb{Q}$}-algebra [Mum70, IV.19, Corollaries 1 and 2].
@@
[…]
Let $k$ be a finite field ${\mathbb{F}}_{q}$. The Frobenius endomorphism $F_{A}:A\to A$ over $k$ is the morphism that is trivial on the topological space and raises functions to their $q$-th powers. The endomorphism $F_{A}$ acts on the Tate module by a semi-simple linear transformation, which we denote by $F:T_{\ell}(A)\to T_{\ell}(A)$. The characteristic polynomial
@@
is called the Weil polynomial of AA. It is a monic polynomial of degree $2\dim A$ with rational integer coefficients independent of the choice of $\ell$ (see [Mum70, Page 180, Theorem 4], and [Mum70, Page 205]).
The Frobenius endomorphism generates the center $K_{A}\subset\operatorname{End}^{\circ}(A)$. Tate proved that the isogeny class of an abelian variety over $k$ is determined by its characteristic polynomial, that is $f_{A}(t)=f_{B}(t)$ if and only if AA is isogenous to $B$ [Ta66]. Moreover, for any $\ell\neq p$ we have
@@
[…]
Assume that AA is simple. In this case $K_{A}$ is a field. Let $L_{A}$ be the Galois envelope of $K_{A}$ over $\mathbb{Q}$}, and let $P$ be the set of primes of $L_{A}$ over pp. The slopes of AA are defined as the set
@@
where $v_{\mathfrak{p}}$ is the normalized valuation. We say that AA is supersingular if all slopes of AA are equal to $1/2$.
@@
Let $L$ be a Galois extension of $\mathbb{Q}$}. Assume that $\sqrt{q}\not\in{\mathbb{Z}}$, and $p$ is unramified in $L$. Then there is no supersingular abelian variety AA such that $L_{A}=L$.
@@
Let AA be an abelian variety such that $L_{A}=L$. Then the slopes of AA are the fractions ${v_{\mathfrak{p}}(F_{A})}/{v_{\mathfrak{p}}(q)}$ with odd denominator. ∎
The abelian variety is ordinary if the set of slopes of AA is $\{0,1\}$. It is known that AA is ordinary if and only if the order of the group $A[p]({\bar{k}})$ is equal to $p^{\dim A}$. A simple abelian surface AA is called mixed if the set of slopes of AA is $\{0,1/2,1\}$.
In dimensions 11 and 22 Weil polynomials can be explicitly classified. We present this classification in part. For more details, see [MN02] and [Wa69].
@@
[Wa69, Ru90, MN02] Let AA be a simple abelian surface over ${\mathbb{F}}_{q}$, where $q=p^{n}$. Then $f_{A}(t)=P_{A}(t)^{e}$, where $P_{A}\in{\mathbb{Z}}[t]$ is irreducible.
@@
    If AA is ordinary or mixed, then $e=1$, and
@@
    If AA is supersingular, then we have the following possibilities:
@@
        (0,q)(0,-q), $n$ is even, and $p\not\equiv 1\bmod 12$;
@@
        $(\pm\sqrt{q},q)$, $n$ is even, and $p\not\equiv 1\bmod 5$;